Obtain an expression for the radial velocity of objects in the galaxy as a function of their
galactic longitude.
1
Expert's answer
2019-04-30T09:58:27-0400

Radial velocity components projected on to the line of sight:


"v_r=\theta v_0 \cos{l} – Δv \sin{l} (1)"

Using (1) we got:


"v_r = \frac {v_0} {R_0} r \sin{l} \cos{l} - \frac {dv} {dR} r \sin{l} \cos{l} (2)"

"v_r = \frac {v_0} {R_0} r \sin{l} \cos{l} - \frac {dv} {dR} r \sin{l} \cos{l} (3)"

where


"A = \frac {1} {2} (\frac {v_0} {R_0} - \frac {dv} {dR})"

 is Oort's first (sheer) constant